What are the steps in simplyfiying 2(5+4y)/5 +10y=2 ?
Found 2 solutions by Maths68, stanbon:
Answer by Maths68(1474)

2(5+4y)/5 +10y=2
Multiply by LCM (which is 5) both sides
5*(2(5+4y)/5 +10y)=2*5
2(5+4y)+50y=10
10+8y+50y=10
58y=10-10
58y=0
58y/58=0/58
y=0
Check
2(5+4y)/5 +10y=2
2(5+4(0))/5 +10(0)=2
2(5)/5=2
2=2

Answer by stanbon(75887)

2(5+4y)/5 +10y=2
---------
[(10+8y)/5] + 10y = 2
----
Multiply thru by 5 to get:
10+8y + 50y = 10
Subtract 10 from each side to get:
8y =-50y
That is a contradiction.
---
Solution: There is no solution.
=============================
Cheers,
Stan H.
==========